OpenStudy (abbycross167):

Someone please help show me how to do this? I really need help... I will give a medal and will become a fan~~~ Graph the following equations using the point-slope method or intercepts method. (1 pt) Show the steps of the procedure as outlined in the lesson as you graph the lines. (1 pt) Draw your y-intercept in yellow, your other points in red and your line in blue. 3x - 2y= 18

Vocaloid (vocaloid):

hint: find the intercepts to find the y-intercept, find y when x = 0 3x - 2y= 18 if x = 0, what is y?

OpenStudy (abbycross167):

-2?

Vocaloid (vocaloid):

nope if x = 0, what does 3x = ?

OpenStudy (abbycross167):

0

Vocaloid (vocaloid):

good so 3x - 2y = 18 becomes 0 - 2y = 18 so what is y?

OpenStudy (abbycross167):

umm I don't know

OpenStudy (dariusx):

i kind of explained this to you today, haven't i ? http://openstudy.com//updates/5665d7fbe4b02fbc766c9c6e

Vocaloid (vocaloid):

-2y = 18... divide both sides by 2, what do you get?

OpenStudy (abbycross167):

@DariusX Yes you did and I tried to do this on my own but I couldn't figure it out so I came here for help -2/2=-1 18/2=9

OpenStudy (dariusx):

okay, i hope someone can explain it to you better than me, good luck

OpenStudy (abbycross167):

You did a great job but this problem confused me

OpenStudy (abbycross167):

@Vocaloid are you going to help me? :/

Vocaloid (vocaloid):

oh whoops sorry

Vocaloid (vocaloid):

so -y = 9, meaning that y = -9

Vocaloid (vocaloid):

now we find the x-intercept by setting y = 0 and finding x 3x - 2y = 18 when y = 0, what does x = ?

OpenStudy (abbycross167):

3?

OpenStudy (abbycross167):

@Vocaloid

Vocaloid (vocaloid):

3x = 18 what does x =?

OpenStudy (abbycross167):

6? @Vocaloid My WiFi is really slow so I didn't see your post until just now sorry